# Odd coloring graphs with linear neighborhood complexity

### v1

@@ 3 Odd coloring and neighborhood complexity
In this section we prove Theorem 1.1, which says that bipartite graphs of linear neighborhood complexity have bounded odd chromatic number.
@@
Let $c=c(r)$ be the constant from Haussler’s Shallow Packing Lemma, which is stated as Lemma 2.1. Then set $\delta\coloneqq r\ell c+1$.
Going for a contradiction, suppose there are no two vertices in $Y^{\prime}$ which are $\delta$-near-twins in $G$. Let =$\mathcal{F}=\{N(v):v\in Y^{\prime}\}$, so that $\mathcal{F}$ is a collection of subsets of $X$. Since $Y^{\prime}$ does not contain any pair of twins, $|\mathcal{F}|=|Y^{\prime}|$. Moreover, since $Y^{\prime}$ does not contain any pair of $\delta$-near-twins, $\mathcal{F}$ is $\delta$-separated. Finally, by the assumption about the neighborhood complexity of $G$, we have that $\pi_{\mathcal{F}}(m)\leqslant rm$ for every positive $m$. So $|Y^{\prime}|\leqslant c|X|/\delta$ by Lemma 2.1. However, since $X$ contains no twins, we also have
@@
Now we are ready to prove Theorem 1.1. We restate the theorem below for convenience.
See 1.1
@@
Let $(X,Y)$ be a bipartition of $G$. We claim that it suffices to show how to color $Y$ so that every non-isolated vertex in $X$ has a color that appears an odd number of times in its neighborhood. Then, if $\varphi_{Y}:Y\rightarrow\mathbb{N}$ is the coloring of $Y$, and $\varphi_{X}:X\rightarrow\mathbb{N}$ is the coloring of $X$ (which exists by symmetry between $X$ and $Y$), we can take the final coloring to be
$\phi(v)=\begin{cases}(1,\varphi_{X}(v))&\textrm{if }v\in X\cr(2,\varphi_{Y}(v))&\textrm{if }v\in Y.\end{cases}$
Taking this product coloring ensures that $\phi$ is proper.
Let $\delta$ be the function from Lemma 3.1, and set $k\coloneqq\max\left(2r+1,\delta(r,2r+1)+1\right)$. We now show that such a coloring $\varphi_{Y}$ exists using at most $k$ colors. Going for a contradiction, suppose otherwise. Let $G$ be a graph with a bipartition $(X,Y)$ which forms a counterexample. Choose $G,X,Y$ so that $|V(G)|$ is minimum. First, we prove two claims.
@@
Our goal is to apply Lemma 3.1 to find a pair of near-twins in $Y$. Then we will use the same color for those two vertices, and apply the fact that $G$ is a minimum counterexample. However, we will have to avoid pairs of vertices $u,v\in Y$ so that there exists a degree-$2$ vertex in $X$ whose neighbors are $u$ and $v$. (Such a vertex forces us to use different colors for $u$ and $v$.) The next claim gives us a large subset of $Y$ to work with.
@@
Let $Y^{\prime}\subseteq Y$ be as in Claim 2. Notice that $|Y^{\prime}|\geqslant 2$ since otherwise $|Y|\leqslant 2r+1$ and we could color every vertex in $Y$ with a different color. Now, notice that $G,X,Y,Y^{\prime}$ satisfy the conditions of Lemma 3.1 with $\ell\coloneqq 2r+1$. (Recall that $X$ contains no pair of twins by Claim 1.) Thus, there exist distinct vertices $u,v\in Y^{\prime}$ which are $(k-1)$-near-twins. By the minimality of $G$, there exists a coloring $\varphi:Y\setminus\{u,v\}\rightarrow\mathbb{N}$ with at most $k$ colors so that every vertex in $X$ with at least one neighbor in $Y\setminus\{u,v\}$ has a color that appears an odd number of times in its neighborhood. By the choice of $Y^{\prime}$, this is actually every vertex in $X$ except for possibly some vertices of degree 00 or $1$. We do not need to worry about the vertices in $X$ of degree 00 or $1$ as they automatically satisfy the condition.
@@
4 Applications
In this section we prove the two corollaries of Theorem 1.1 (Corollary 1.2 and Corollary 1.3).
We begin by using the Growth Rate Theorem (stated as Theorem 2.4) to control the neighborhood complexity of certain fundamental graphs. We note that we could do a “one-sided” version of the following lemma. However this version, which is symmetric under taking duals, is more convenient for our purposes. Note that Corollary 1.3 immediately follows from Theorem 1.1 and the following lemma.
Lemma 4.1.
For any integers $t$ and $q$, there exists an integer $r=r(t,q)$ so that if $M$ is a matroid with no $U_{2,q+2}$, $U_{q,q+2}$, $M(K_{t})$, or $M(K_{t})^{*}$ minor, and $B$ is a basis of $M$, then the neighborhood complexity of $\mathcal{F}(M,B)$ is at most $r\cdot m$ for every positive $m$.
Proof.
Let $g$ denote the function from Theorem 2.4; so any simple rank-$n$ matroid with no $U_{2,q+2}$ or $M(K_{t})$ minor has at most $g(t,q)\cdot n$ elements. Let $M$ be a matroid with no $U_{2,q+2}$, $U_{q,q+2}$, $M(K_{t})$, or $M(K_{t})^{*}$ minor, and let $B$ be a basis of $M$.
Since $\mathcal{F}(M,B)$ is bipartite, and a matroid and its dual have the same fundamental graph by Lemma 2.2, it suffices to consider the number of different neighborhoods within a set $A\subseteq B$. Let $F\subseteq E(M)\setminus B$ be obtained by selecting one vertex $v$ with $N(v)\cap A=A^{\prime}$ for each subset $A^{\prime}\subseteq A$ for which $|A^{\prime}|\geqslant 2$, and such a vertex $v$ exists. Thus |$|\{N(v)\cap A:v\in V(\mathcal{F}(M,B))\}|\leqslant|F|+|A|+1$ because $A$ has $|A|+1$ subsets with at most one element.
Now, let $M^{\prime}$ be the matroid obtained from $M$ by contracting all elements in $B\setminus A$ and deleting all elements in $(E(M)\setminus B)\setminus F$. By Lemma 2.3, the subgraph of $\mathcal{F}(M,B)$ induced on $A\cup F$ is equal to $\mathcal{F}(M^{\prime},A)$. Note that $M^{\prime}$ is a rank-$|A|$ matroid with no $U_{2,q+2}$ or $M(K_{t})$ minor. We claim that $M^{\prime}$ is simple as well. This will complete the proof since then the Growth Rate Theorem implies that $|F|+|A|=|E(M^{\prime})|\leqslant g(t,q)\cdot|A|$.
First of all, $M^{\prime}$ has no loops since $A$ is a basis and every vertex in $F$ has at least one neighbor in $A$. Next, $M^{\prime}$ has no parallel pair with one element in $A$ and the other in $F$ since every element in $F$ has at least two neighbors in $A$. Finally, $M^{\prime}$ has no parallel pair with both elements in $F$ since no two vertices in $F$ have the same neighborhood in $A$. So $M^{\prime}$ is simple, which completes the proof. ∎
Recall that a class of graphs is $\chi_{odd}$-bounded if there exists a function $f$ so that any graph in the class with clique number $\omega$ has odd chromatic number at most $f(\omega)$. Similarly, a class is $\chi$-bounded if there exists a function $g$ so that any graph in the class with clique number $\omega$ has chromatic number at most $g(\omega)$. We denote the chromatic number of a graph $G$ by $\chi(G)$. We say that a class is hereditary if it is closed under taking induced subgraphs.
The following lemma from [29] implies that in order to understand whether a hereditary $\chi$-bounded class is also $\chi_{odd}$-bounded, it suffices to consider the bipartite graphs in the class. We state a slightly weaker but simpler version of the lemma.
Lemma 4.2 ([29, Lemma 1.3]).
If every induced bipartite subgraph of a graph $G$ has odd chromatic number at most $t$, then $\chi_{odd}(G)\leqslant t^{2}\chi(G)$.
Now we may prove Corollary 1.2. We omit the definitions of twin-width, merge-width, and vertex-minors; instead we refer the reader to [6][18], and [13], respectively. The corollary is restated below for convenience.
@@ 4 Applications
We may assume that $\mathcal{G}$ is hereditary because taking induced subgraphs cannot increase the twin-width or merge-width, or create a new vertex-minor. Recall that classes with bounded twin-width [5, 9, 39], classes with bounded merge-width [3], and classes with a forbidden vertex-minor [13] are known to be $\chi$-bounded. Thus, by Lemma 4.2 and Theorem 1.1, it suffices to prove that every bipartite graph $H\in\mathcal{G}$ has linear neighborhood complexity. For classes of bounded twin-width, this was proven by Bonnet, Foucaud, Lehtilä, and Parreau [4]. For classes with bounded merge-width, this was recently proven by Bonamy and Geniet [3].
For classes with a forbidden vertex-minor, we observe that the $1$-subdivision of $K_{t}$ contains every $t$-vertex graph as a vertex-minor, and is itself the fundamental graph of $K_{t+1}$ with respect to the $(t+1)$-vertex star. Since every pivot-minor is also a vertex-minor, it follows that any class of bipartite graphs with a forbidden vertex-minor also forbids the $1$-subdivision of $K_{t}$ as a pivot-minor (for some $t$). Finally, due to the connection between pivot-minors of bipartite graphs and minors of binary matroids [35], it follows from Lemma 4.1 that bipartite graphs with a forbidden vertex-minor have linear neighborhood complexity, as desired. ∎
5 Odd coloring and bipartite circle graphs
@@
Given a graph $G$ and a spanning forest $T$ of $G$, the fundamental cut of an edge $f\in E(T)$ is the set $C^{*}(T,f)$ which contains all edges in $E(G)$ which join the vertex-sets of the two components of $T-f$ which are not components of $T$. Note that $f$ itself is included in $C^{*}(T,f)$. Indeed, $C^{*}(T,f)\setminus\{f\}$ is the neighborhood of $f$ in $\mathcal{F}(G,T)$.
@@
For any graph $G$, any spanning forest $T$ of $G$, and any edge $f_{0}$ of $T$, there exists a coloring $\varphi:C^{*}(T,f_{0})\setminus\{f_{0}\}\rightarrow\{1,\ldots,9\}$ so that for each tree-edge $f\in E(T)$ with $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\neq\emptyset$ there exists a color $i\in\{1,\ldots,8\}$ which occurs an odd number of times on $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})$.
@@
Let $T_{1}$ and $T_{2}$ be the two components of $T-f_{0}$. We shall color the non-tree edges in the fundamental cut $C^{*}(T,f_{0})$ of $f_{0}$, i.e., the non-tree edges with one endpoint in $T_{1}$ and the other endpoint in $T_{2}$. Consider the tree $T_{1}^{\prime}=T_{1}+f_{0}$ to be rooted at the endpoint of $f_{0}$ in $T_{2}$. Observe that for any edge $f$ in $T_{1}^{\prime}$, a non-tree edge $e\in C^{*}(T,f_{0})$ lies in $C^{*}(T,f)$ if and only if the endpoint of $e$ in $T_{1}$ lies in the subtree of $T_{1}^{\prime}$ below $f$. It follows that the set family $\mathcal{S}_{1}=\{C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\colon f\in T_{1}^{\prime}\}$ is laminar, i.e., its members are either disjoint or in a containment relation. From [29, Lemma 3.8] we get that there is a coloring $\varphi_{1}:C^{*}(T,f_{0})\setminus\{f_{0}\}\to\{1,2,3\}$ such that every element of $\mathcal{S}_{1}$ has an odd number of elements colored $1$ or an odd number of elements colored $2$.
Symmetrically, there is a coloring $\varphi_{2}:C^{*}(T,f_{0})\setminus\{f_{0}\}\to\{1,2,3\}$ such that every element of $\mathcal{S}_{2}=\{C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\colon f\in T_{2}^{\prime}\}$ has an odd number of elements colored $1$ or an odd number of elements colored $2$. Setting $\varphi(e)=(\varphi_{1}(e),\varphi_{2}(e))$ for every $e\in C^{*}(T,f_{0})\setminus\{f_{0}\}$, we obtain the desired $9$-coloring. Indeed, if $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})$ has an odd number of elements colored $i$ in $\varphi_{1}$, then in $\varphi$ these elements are distributed into three color classes $(i,1)$, $(i,2)$, $(i,3)$, at least one of which is necessarily odd. Note that since by [29, Lemma 3.8] in both factors of the product coloring the color $3$ is not needed in order to witness having an odd color in the neighborhoods, we can designate color $(3,3)$ as color $9$ from the statement. ∎
For the next result, recall that the star-arboricity $\mathrm{sa}(G)$ of a graph $G$ is the smallest $\ell$, such that $E(G)$ can be decomposed into $\ell$ star forests, i.e., each component has diameter at most $2$. Note that the $\mathrm{sa}(G)$ is bounded by twice the degeneracy of $G$, see [1]. Hence, in particular for any integer $t$, there exists an integer $\ell(t)$ so that if $G$ is a graph with no $K_{t}$ minor, then $\mathrm{sa}(G)\leqslant\ell(t)$. Indeed, by the Kostochka–Thomason bound $\ell(t)\in O(t\sqrt{\log t})$, see [31, 32, 42, 41].
Theorem 5.2.
For any integer $t$, there exists an integer $k=k(t)$ so that if $G$ is a graph with no $K_{t}$ minor, and $T$ is a spanning forest of $G$, then there exists a coloring $\phi:E(G)\setminus E(T)\rightarrow\{1,\ldots,k\}$ so that for each non-bridge edge $f\in E(T)$, there exists a color $i\in\{1,\ldots,k\}$ which occurs an odd number of times on $C^{*}(T,f)$. Indeed, $k(t)\leqslant 16\ell(t-1)+1$, with $\ell(\cdot)$ defined as above the theorem.
@@
We root $T$ at an arbitrary vertex $r$. We shall construct a sequence $G_{0}\subseteq G_{1}\subseteq\cdots\subseteq G_{x}=G$ of subgraphs of $G$, starting with $G_{0}$ being just vertex $r$. For a non-tree edge $e$, the fundamental cycle $C(T,e)$ is the unique cycle in $T+e$. Note that for a tree-edge $f$ and a non-tree edge $e$ we have $f\in C(T,e)$ if and only if $e\in C^{*}(T,f)$. For $i>0$, having defined $G_{i-1}$, let
@@
Next, we color the edges in $B_{i}$ (which is non-empty only for $i>0$), i.e., the non-tree edges in $G_{i}$ that are not in $G_{i-1}$. To this end, consider $G_{i}$ with spanning tree $T_{i}$. Contract the subtree $T_{i-1}$ of $T_{i}$ to a single vertex $a$. Also, consider the components of $T_{i}$ after the removal of all vertices in $T_{i-1}$ and contract each such component to a single vertex. The resulting graph $H$ is a minor of $G_{i}$ with two vertices in $H-a$ being adjacent if and only if some non-tree edge in $B_{i}$ connects the corresponding subtrees of $T_{i}-V(T_{i-1})$.
Since $a$ is a universal vertex in $H$, the graph $H-a$ has no $K_{t-1}$ minor. Thus, as argued above the theorem there exists a universal constant $\ell=\ell(t-1)$ such that $E(H-a)$ can be decomposed into $\ell$ star forests $F_{1},\ldots,F_{\ell}$. (We may assume that each vertex of $H-a$ is the center of at least one star.) For each star $S$ in $F_{1},\ldots,F_{\ell}$, consider its center $v$ in $H-a$, the corresponding component $K$ of $T_{i}-V(T_{i-1})$, and the edge $t_{0}\in T_{i}$ that connects $K$ and $T_{i-1}$. Observe that $C^{*}(T,f_{0})\setminus\{t_{0}\}\subseteq B_{i}$. Let $B(S)\subseteq B_{i}$ be the subset of non-tree edges that correspond in $H$ to edges in the star $S$ or the edge between $a$ and $v$. By Lemma 5.1 there is a $9$-coloring $\varphi_{S}$ of $B(S)$ such that for every $f\in T$ the set $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\cap B(S)$ is either empty or contains a color an odd number of times.

### v2

@@ 3 Odd coloring and neighborhood complexity
In this section we prove Theorem˜1.2, which says that graphs of linear neighborhood complexity have bounded improper odd chromatic number.
@@
Let $c=c(r)$ be the constant from Haussler’s Shallow Packing Lemma, which is stated as Lemma˜2.1. Then set $\delta\coloneqq r\ell c+1$.
Going for a contradiction, suppose there are no two vertices in $Y^{\prime}$ which are $\delta$-near-twins in $G$. Let $\mathcal{F}=\{N(v):v\in Y^{\prime}\}$, so that $\mathcal{F}$ is a collection of subsets of $X$. Since $Y^{\prime}$ does not contain any pair of twins, $|\mathcal{F}|=|Y^{\prime}|$. Moreover, since $Y^{\prime}$ does not contain any pair of $\delta$-near-twins, $\mathcal{F}$ is $\delta$-separated. Finally, by the assumption about the neighborhood complexity of $G$, we have that $\pi_{\mathcal{F}}(m)\leqslant rm$ for every positive $m$. So $|Y^{\prime}|\leqslant c|X|/\delta$ by Lemma˜2.1. However, since $X$ contains no twins, we also have
@@
We are ready to prove a one-sided odd-coloring result for bipartite graphs, which will be the main ingredient to the full theorem.
Lemma 3.2.
For any integer $r$, there exists an integer $k=k(r)$ so that if $G$ is a bipartite graph with bipartition $(X,Y)$ and neighborhood complexity $\eta_{G}(m)\leqslant r\cdot m$ for every positive $m$, then $Y$ can be $k$-colored such that for each non-isolated vertex $v\in X$, there exists a color which appears an odd number of times in $N(v)$.
@@
Let $\delta$ be the function from Lemma˜3.1, and set $k\coloneqq\max\left(2r+1,\delta(r,2r+1)+1\right)$. We now show that a coloring $\varphi_{Y}$ as claimed exists using at most $k$ colors. Going for a contradiction, suppose otherwise. Let $G$ be a graph with a bipartition $(X,Y)$ which forms a counterexample. Choose $G,X,Y$ so that $|V(G)|$ is minimum. First, we prove two claims.
@@
Our goal is to apply Lemma˜3.1 to find a pair of near-twins in $Y$. Then we will use the same color for those two vertices, and apply the fact that $G$ is a minimum counterexample. However, we will have to avoid pairs of vertices $u,v\in Y$ so that there exists a degree-$2$ vertex in $X$ whose neighbors are $u$ and $v$. (Such a vertex forces us to use different colors for $u$ and $v$.) The next claim gives us a large subset of $Y$ to work with.
@@
Let $Y^{\prime}\subseteq Y$ be as in Claim˜2. Notice that $|Y^{\prime}|\geqslant 2$ since otherwise $|Y|\leqslant 2r+1$ and we could color every vertex in $Y$ with a different color. Now, notice that $G,X,Y,Y^{\prime}$ satisfy the conditions of Lemma˜3.1 with $\ell\coloneqq 2r+1$. (Recall that $X$ contains no pair of twins by Claim˜1.) Thus, there exist distinct vertices $u,v\in Y^{\prime}$ which are (k1)(k-1)-near-twins. By the minimality of $G$, there exists a coloring $\varphi:Y\setminus\{u,v\}\rightarrow\mathbb{N}$ with at most $k$ colors so that every vertex in $X$ with at least one neighbor in $Y\setminus\{u,v\}$ has a color that appears an odd number of times in its neighborhood. By the choice of $Y^{\prime}$, this is actually every vertex in $X$ except for possibly some vertices of degree 0 or $1$. We do not need to worry about the vertices in $X$ of degree 0 or $1$ as they automatically satisfy the condition.
@@
Now we are ready to prove Theorem˜1.2. We restate the theorem below for convenience.
@@
Denote the tensor product $H=G\times K_{2}$, with vertex set $V(H)=V(G)\times[2]$ and an edge $\{(u,i),(v,j)\}$ if and only if $\{u,v\}\in E(G)$ and $\{i,j\}=[2]$. If $A\subseteq V(G)\times[2]$ denote $A_{i}=\{v\in V(G)\mid(v,i)\in A\}$ for $i=1,2$. Then for any $A\subseteq V(G)\times[2]$,
$|\{N(v,i)\cap A:(v,i)\in V(H)\}|\leqslant|\{N(v)\cap A_{1}:v\in V(G)\}|+|\{N(v)\cap A_{2}:v\in V(G)\}|.$
If $|A|=m$, then the latter is at most $2\eta_{G}(m)$. Hence $\eta_{H}(m)\leqslant 2rm$ for all $m$. Further, $H$ is bipartite with bipartition $(X,Y)=(V(G)\times\{1\},V(G)\times\{2\})$. Thus, we can color $Y$ with $k(2r)$ colors so that every non-isolated vertex in $X$ has a color that appears an odd number of times in its neighborhood, where $k$ is the function from Lemma˜3.2 . Since $Y=V(G)\times\{2\}$, this coloring can be interpreted as an improper odd coloring of $G$. ∎
4 Applications
In this section we prove the two corollaries of Theorem˜1.2 (Corollary˜1.3 and Corollary˜1.4).
We begin by using the Growth Rate Theorem (stated as Theorem˜2.4) to control the neighborhood complexity of certain fundamental graphs. We note that we could do a “one-sided” version of the following lemma. However this version, which is symmetric under taking duals, is more convenient for our purposes. Note that Corollary˜1.4 immediately follows from Theorem˜1.2 and the following lemma.
Lemma 4.1.
For any integers $t$ and $q$, there exists an integer $r=r(t,q)$ so that if $M$ is a matroid with no $U_{2,q+2}$, $U_{q,q+2}$, $M(K_{t})$, or $M(K_{t})^{*}$ minor, and $B$ is a basis of $M$, then the neighborhood complexity $\eta_{\mathcal{F}(M,B)}(m)$ of $\mathcal{F}(M,B)$ is at most $r\cdot m$ for every positive $m$.
Proof.
Let $g$ denote the function from Theorem˜2.4; so any simple rank-$n$ matroid with no $U_{2,q+2}$ or $M(K_{t})$ minor has at most $g(t,q)\cdot n$ elements. Let $M$ be a matroid with no $U_{2,q+2}$, $U_{q,q+2}$, $M(K_{t})$, or $M(K_{t})^{*}$ minor, and let $B$ be a basis of $M$.
Since $\mathcal{F}(M,B)$ is bipartite, and a matroid and its dual have the same fundamental graph by Lemma˜2.2, it suffices to consider the number of different neighborhoods within a set $A\subseteq B$. Let $F\subseteq E(M)\setminus B$ be obtained by selecting one vertex $v$ with $N(v)\cap A=A^{\prime}$ for each subset $A^{\prime}\subseteq A$ for which $|A^{\prime}|\geqslant 2$, and such a vertex $v$ exists. Thus $|\{N(v)\cap A:v\in V(\mathcal{F}(M,B))\}|\leqslant|F|+|A|+1$ because $A$ has $|A|+1$ subsets with at most one element.
Now, let $M^{\prime}$ be the matroid obtained from $M$ by contracting all elements in $B\setminus A$ and deleting all elements in $(E(M)\setminus B)\setminus F$. By Lemma˜2.3, the subgraph of $\mathcal{F}(M,B)$ induced by $A\cup F$ is equal to $\mathcal{F}(M^{\prime},A)$. Note that $M^{\prime}$ is a rank-$|A|$ matroid with no $U_{2,q+2}$ or $M(K_{t})$ minor.
We claim that $M^{\prime}$ is simple as well. First of all, $M^{\prime}$ has no loops since $A$ is a basis and every vertex in $F$ has at least one neighbor in $A$. Next, $M^{\prime}$ has no parallel pair with one element in $A$ and the other in $F$ since every element in $F$ has at least two neighbors in $A$. Finally, $M^{\prime}$ has no parallel pair with both elements in $F$ since no two vertices in $F$ have the same neighborhood in $A$. So $M^{\prime}$ is simple, and the Growth Rate Theorem implies that $|F|+|A|=|E(M^{\prime})|\leqslant g(t,q)\cdot|A|$, which completes the proof. ∎
Recall that a class of graphs is $\chi_{\mathrm{o}}$-bounded if there exists a function $f$ so that any graph in the class with clique number $\omega$ has odd chromatic number at most $f(\omega)$. Similarly, a class is $\chi$-bounded if there exists a function $g$ so that any graph in the class with clique number $\omega$ has chromatic number at most $g(\omega)$. We denote the chromatic number of a graph $G$ by $\chi(G)$. We say that a class is hereditary if it is closed under taking induced subgraphs.
The following lemma from [37] implies that in order to understand whether a hereditary $\chi$-bounded class is also $\chi_{\mathrm{o}}$-bounded, it suffices to consider the bipartite graphs in the class. We state a slightly weaker but simpler version of the lemma.
Lemma 4.2 ([37, Lemma 1.3]).
If every induced bipartite subgraph of a graph $G$ has odd chromatic number at most $t$, then $\chi_{\mathrm{o}}(G)\leqslant t^{2}\chi(G)$.
Now we may prove Corollary˜1.3. We omit the definitions of merge-width and vertex-minors; instead we refer the reader to [22] and [17], respectively. The corollary is restated below for convenience.
See 1.3
Proof.
First suppose that $\mathcal{G}$ has bounded merge-width. Bonamy and Geniet [2] recently proved that classes with bounded merge-width are $\chi$-bounded and have linear neighborhood complexity. Thus, by Remark˜1.1 and Theorem˜1.2, the class $\mathcal{G}$ is $\chi_{\mathrm{o}}$-bounded.
Now suppose that $\mathcal{G}$ has a forbidden vertex-minor. We may assume that $\mathcal{G}$ is hereditary because taking induced subgraphs cannot create a new vertex-minor. Moreover, $\mathcal{G}$ is $\chi$-bounded by [17]. Thus, by Lemma˜4.2, it suffices to prove that the bipartite graphs in $\mathcal{G}$ have bounded odd chromatic number.
We observe that the $1$-subdivision of $K_{t}$ contains every $t$-vertex graph as a vertex-minor. Since every pivot-minor is also a vertex-minor, it follows that any class of bipartite graphs with a forbidden vertex-minor also forbids the $1$-subdivision of $K_{t}$ as a pivot-minor (for some $t$). Now, consider a bipartite graph $G\in\mathcal{G}$, and let $A$ denote the bipartite adjacency matrix of $G$. (So the rows of $A$ correspond to one side of the bipartition, the columns of $A$ correspond to the other, and each entry is 0 or $1$ depending on whether the corresponding vertices are adjacent.) Let $M$ denote the column matroid of the binary matrix $I\text{ | }A]$ where $I$ is an identity matrix whose size equals the number of rows of $A$. Then $G$ is the fundamental graph of $M$ with respect to the basis corresponding to the columns of $I$. Also notice that the $1$-subdivision of $K_{t}$ is itself the fundamental graph of $M(K_{t+1})$ with respect to the basis formed by the edges of a $(t+1)$-vertex star.
Bouchet [6] proved a connection between minors of binary matroids and pivot-minors of their fundamental graphs which yields the following result; if $M(K_{t+1})$ or its dual is a minor of $M$, then the $1$-subdivision of $K_{t}$ is a pivot-minor of $G$. It follows that $M$ forbids both $M(K_{t+1})$ and its dual as minors. Since $M$ is binary, it also forbids $U_{2,4}$ as a minor. So by Lemma˜4.1, the bipartite graphs in $\mathcal{G}$ have linear neighborhood complexity. Thus, by Theorem˜1.2, the bipartite graphs in $\mathcal{G}$ have bounded improper odd chromatic number. This also bounds their odd chromatic number by Remark˜1.1, as desired. ∎
5 Odd coloring fundamental graphs of graphs excluding a minor
@@
Given a graph $G$ and a spanning forest $T$ of $G$, we denote the fundamental cut of $T$ with respect to an edge $f\in E(T)$ by $C^{*}(T,f)$, i.e., it contains all edges in $E(G)$ which join the vertex-sets of the two components of TfT-f which are not components of $T$. Note that $f$ itself is included in $C^{*}(T,f)$. Indeed, $C^{*}(T,f)\setminus\{f\}$ is the neighborhood of $f$ in $\mathcal{F}(G,T)$.
@@
For any graph $G$, any spanning forest $T$ of $G$, and any edge $f_{0}$ of $T$, there exists a coloring $\varphi:C^{*}(T,f_{0})\setminus\{f_{0}\}\rightarrow\{1,\ldots,9\}$ so that for each forest-edge $f\in E(T)$ with $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\neq\emptyset$ there exists a color $i\in\{1,\ldots,8\}$ which occurs an odd number of times on $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})$.
@@
The claimed coloring can then be defined independently on each connected component of $G$. Hence, we assume that $G$ is connected and $T$ is a spanning tree. Let $T_{1}$ and $T_{2}$ be the two components of $T-f_{0}$. We shall color the non-tree edges in the fundamental cut $C^{*}(T,f_{0})$ of $f_{0}$, i.e., the non-tree edges with one endpoint in $T_{1}$ and the other endpoint in $T_{2}$. Consider the tree $T_{1}^{\prime}=T_{1}+f_{0}$ to be rooted at the endpoint of $f_{0}$ in $T_{2}$. Observe that for any edge $f$ in $T_{1}^{\prime}$, a non-tree edge $e\in C^{*}(T,f_{0})$ lies in $C^{*}(T,f)$ if and only if the endpoint of $e$ in $T_{1}$ lies in the subtree of $T_{1}^{\prime}$ below $f$. It follows that the set family $\mathcal{S}_{1}=\{C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\colon f\in T_{1}^{\prime}\}$ is laminar, i.e., its members are either disjoint or in a containment relation. From [37, Lemma 3.8] we get that there is a coloring $\varphi_{1}:C^{*}(T,f_{0})\setminus\{f_{0}\}\to\{1,2,3\}$ such that every element of $\mathcal{S}_{1}$ has an odd number of elements colored $1$ or an odd number of elements colored $2$.
Symmetrically, there is a coloring $\varphi_{2}:C^{*}(T,f_{0})\setminus\{f_{0}\}\to\{1,2,3\}$ such that every element of $\mathcal{S}_{2}=\{C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\colon f\in T_{2}^{\prime}\}$ has an odd number of elements colored $1$ or an odd number of elements colored $2$. Setting $\varphi(e)=(\varphi_{1}(e),\varphi_{2}(e))$ for every $e\in C^{*}(T,f_{0})\setminus\{f_{0}\}$, we obtain the desired $9$-coloring. Indeed, if $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})$ has an odd number of elements colored $i$ in $\varphi_{1}$, then in $\varphi$ these elements are distributed into three color classes $(i,1)$, $(i,2)$, $(i,3)$, at least one of which is necessarily odd. Note that since by [37, Lemma 3.8] in both factors of the product coloring the color $3$ is not needed in order to witness having an odd color in the neighborhoods, we can designate color $(3,3)$ as color $9$ from the statement. ∎
We are ready to prove the main theorem of the section, where we recall that $\ell(t)$ is the smallest integer so that if $G$ is a graph with no $K_{t}$ minor, then the star-arboricity of $G$ is at most $\ell(t)$.
See 1.5
@@
We root $T$ at an arbitrary vertex $r$. We shall construct a sequence $G_{0}\subseteq G_{1}\subseteq\cdots\subseteq G_{x}=G$ of subgraphs of $G$, starting with $G_{0}$ being just vertex $r$. For a non-tree edge $e$, the fundamental cycle $C(T,e)$ is the unique cycle in $T+e$ and for a tree-edge $f$ denote by $C^{*}(T,f)$ the fundamental cut of $T$ with respect to $f$. Note that for a tree-edge $f$ and a non-tree edge $e$ we have $f\in C(T,e)$ if and only if $e\in C^{*}(T,f)$. For $i>0$, having defined $G_{i-1}$, let
@@
Next, we color the edges in $B_{i}$ (which is non-empty only for $i>0$), i.e., the non-tree edges in $G_{i}$ that are not in $G_{i-1}$. To this end, consider $G_{i}$ with spanning tree $T_{i}$. Contract the subtree $T_{i-1}$ of $T_{i}$ to a single vertex $a$. Also, consider the components of $T_{i}$ after the removal of all vertices in $T_{i-1}$ and contract each such component to a single vertex. The resulting graph $H$ is a minor of $G_{i}$ with two vertices in HaH-a being adjacent if and only if some non-tree edge in $B_{i}$ connects the corresponding subtrees of $T_{i}-V(T_{i-1})$.
Since $a$ is a universal vertex in $H$, the graph HaH-a has no $K_{t-1}$ minor. Thus, as argued above the theorem there exists a universal constant $\ell=\ell(t-1)$ such that E(Ha)E(H-a) can be decomposed into $\ell$ star forests $F_{1},\ldots,F_{\ell}$. (We may assume that each vertex of HaH-a is the center of at least one star.) For each star $S$ in $F_{1},\ldots,F_{\ell}$, consider its center $v$ in HaH-a, the corresponding component $K$ of $T_{i}-V(T_{i-1})$, and the edge $t_{0}\in T_{i}$ that connects $K$ and $T_{i-1}$. Observe that $C^{*}(T,f_{0})\setminus\{t_{0}\}\subseteq B_{i}$. Let $B(S)\subseteq B_{i}$ be the subset of non-tree edges that correspond in $H$ to edges in the star $S$ or the edge between $a$ and $v$. By Lemma˜5.1 there is a $9$-coloring $\varphi_{S}$ of $B(S)$ such that for every $f\in T$ the set $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\cap B(S)$ is either empty or contains a color an odd number of times.